mary is three times as old as her son .after twelvue years she is one year less than twice her sons age. what are there present ages.?
Found 2 solutions by scott8148, drk:
Answer by scott8148(6628)

m = 3s

m + 12 = 2(s + 12) - 1 ___ m + 12 = 2s + 24 -1 ___ m = 2s + 11

3s = 2s + 11 ___ s = 11

m = 33

Answer by drk(1908)

Let m = mary, and s = son.
From above, we get
m = 3s
m+12 = 2(s+12) - 1
By substitution, we get
3s + 12 = 2s + 23
s = 11
m = 33